OpenStudy (anonymous):

An equation is shown below: 4(x - 3) - 5(x + 1) = 3 Which statement shows a correct next step in solving the equation? 1) The equation can become 4x - 3 - 5x + 1 = 3 by applying the associative property of multiplication. 2) The equation can become 4x - 3 - 5x + 1 = 3 by applying the distributive property. 3) The equation can become 4x - 12 - 5x - 5 = 3 by applying the distributive property. 4) The equation can become 4x - 12 - 5x - 5 = 3 by applying the associative property of multiplication.

OpenStudy (imstuck):

3

OpenStudy (anonymous):

do you need an explanation?

OpenStudy (anonymous):

Yeah, is the answer 3 though?

OpenStudy (anonymous):

well the answer is 3 because well you have to distribute you know that. then you will get 4x - 12 - 5x - 5 = 3. Distributive property: a(b + c) = ab + ac Associative property: a + (b + c) = (a + b) + c

OpenStudy (anonymous):

4(x - 3) - 5(x + 1) = 3 4x -12 - 5x -5 = 3

OpenStudy (anonymous):

Oh, I see. Thank you :)